# 3d Topological Field Theories and Nahm Sum Formulas

### v2

@@ 1 Introduction
where each of the $U(1)$ factors has precisely one charged chiral multiplet of charge 1.  [^2] See recent works Kucharski:2017ogk ; Ekholm:2020lqy ; Jockers:2021omw ; Chung:2023qth ; Gang:2023rei ; Gang:2023ggt ; Ferrari:2023fez ; Creutzig:2024ljv ; Gaiotto:2024ioj , which focuses on a similar observation. In this context, the $r\times r$ matrix $A$ is identified with the mixed Chern-Simons level matrix $K$. As discussed extensively in recent literature Costello:2018fnz ; Costello:2018swh ; Costello:2020ndc ; Creutzig:2021ext ; Coman:2023xcq , the boundary vertex operator algebras of 3d supersymmetric gauge theories are generically non-rational. However, there are two important classes of exceptions:
@@ 3.1 ${\cal N}=2$ abelian Chern-Simons matter theories
The reference R-charge of the monopole operator $R_{*}(V_{\bf m})$ is
@@
In order to identify a candidate theory ${\cal T}[K,\{{\cal O}_{I}\}]$ which flows to an IR ${\cal N}=4$ rank-0 SCFT, we will require that there are (r1)(r-1) 1/2-BPS CPOs in the superpotential $\cal W$}, i.e. $N_{\cal O}=r-1$. This is because an ${\cal N}=4$ rank-0 theory should possess only one (non-R) $U(1)$ flavor symmetry when described in terms of an ${\cal N}=2$ subalgebra.777 However, it is possible for some ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{N_{\cal O}}]$ with $N_{\cal O}<r-1$ flow to an ${\cal N}=4$ rank-0 SCFT, with some UV symmetries decoupling in the IR. We will discuss such exceptional cases in section 4.3. In this case, the affine subspace is one-dimensional and can be represented as
@@ 3.4 Half-index and modular functions
Finally, we consider an ${\cal N}=2$ gauge theory $T[K,\{{\cal O}_{I}\}]$ defined on the Euclidean half-space $\mathbb{C}\times\mathbb{R}_{\leq 0}$. The local operators on the boundary are counted by the half-index, which can be understood as the partition function on the cigar geometry $D^{2}\times S^{1}$, with a choice of boundary condition $\mathbb{B}$ specified on the boundary torus. Gadde:2013sca ; Gadde:2013wq ; Yoshida:2014ssa ; Dimofte:2017tpi Let $(z,\bar{z},t)$ be the local coordinate around a point on the boundary placed at $t=0$. The half index is defined by
@@
where $c\neq 0$ is a complex number. We denote this boundary condition by $D_{c}$, a deformed Dirichlet boundary condition. The half-index of ${\cal T}[K,\{{\cal O}_{I}\}]$ with the insertion of a supersymmetric Wilson loop $W_{\vec{Q}}$ then reads
@@
with the Pochhammer symbol $(q)_{n}:=\prod_{i=1}^{n}(1-q^{i})$ defined in the usual way. The summation over $m$ labels the boundary monopole operators discussed above.
@@ 4.1.1 $r=1$
    $K=(2)=C(A_{1})$, 121212 Taking into account the background CS level 1/2-1/2 in the ${\cal T}_{\Delta}$ theory (2), this theory corresponds to the $U(1)_{3/2}$ theory coupled to a single chiral multiplet of charge +1+1, which we denote simply as $U(1){3/2}+\Phi$. This theory is IR dual to $U(1)_{-3/2}+\Phi$ Gang:2018huc .
@@ A.1 General structure of partition functions on $\mathcal{M}_{g,p}$
The effective twisted superpotential $\mathcal{W}(u,v)$ is broken down into two pieces, one originating from the Chern-Simons terms of the theory and one originating from the one-loop contribution of the $\Phi_{i}$ [^18] Importantly this is all done in the $U(1)_{-\frac{1}{2}}$ quantization” which is arises due to the parity anomaly that exists in 3D gauge theories as discussed in Closset:2018ghr .:

### v3

@@ 1 Introduction
where each of the $U(1)$ factors has precisely one charged chiral multiplet of charge 1, under a particular choice of boundary condition that will be discussed in section 3.4.222See recent works Kucharski:2017ogk ; Ekholm:2020lqy ; Jockers:2021omw ; Chung:2023qth ; Gang:2023rei ; Gang:2023ggt ; Ferrari:2023fez ; Creutzig:2024ljv ; Gaiotto:2024ioj , which focuses on a similar observation. In this context, the $r\times r$ matrix $A$ is identified with the mixed Chern-Simons level matrix $K$. As discussed extensively in recent literature Costello:2018fnz ; Costello:2018swh ; Costello:2020ndc ; Creutzig:2021ext ; Coman:2023xcq , the boundary vertex operator algebras of 3d supersymmetric gauge theories are generically non-rational. However, there are two important classes of exceptions:
@@ 3.1 ${\cal N}=2$ abelian Chern-Simons matter theories
The reference R-charge of the monopole operator $R_{*}(V_{\bf m})$ is  [^7] The statement in Gang:2018huc that $R_{*}(\Phi)=1/3$ at the superconformal point (for $r=1$, K=1K=-1) should be interpreted with care. The R-charge of a gauge-variant operator is not uniquely fixed by F-maximization, and one must choose a reference point. Our choice $R^{\text{here}}_{*}(\Phi)=0$ and the assignement $R^{\text{GY}}_{*}(\Phi)=1/3$ in loc.cit. are related by $R^{\text{GY}}=R^{\text{here}}+G/3$, with $G$ the gauge charge. Correspondingly, we take the mixed gauge-R CS level to be $k^{\rm our}_{GR}=\tfrac{1}{2}$ while $k^{\rm GY}_{GR}=0$. Applying this to eqn. (20), we obtain $R^{\rm here}_{*}(V_{m})=\tfrac{1}{2}\,(m+|m|)$ and $R^{\rm GY}_{*}(V_{m})=\tfrac{1}{3}|m|$ which is consistent with the relation $R^{\rm GY}=R^{\rm here}+G/3$ since $G(V_{m})=-\tfrac{3}{2}m-\tfrac{1}{2}|m|$. Since the superconformal index receives contributions only from gauge invariant operators, the result is independent of this choice.
@@
In order to identify a candidate theory ${\cal T}[K,\{{\cal O}_{I}\}]$ which flows to an IR ${\cal N}=4$ rank-0 SCFT, we will require that there are (r1)(r-1) 1/2-BPS CPOs in the superpotential $\cal W$}, i.e. $N_{\cal O}=r-1$. This is because an ${\cal N}=4$ rank-0 theory should possess only one (non-R) $U(1)$ flavor symmetry when described in terms of an ${\cal N}=2$ subalgebra. [^8] However, it is possible for some ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{N_{\cal O}}]$ with $N_{\cal O}<r-1$ flow to an ${\cal N}=4$ rank-0 SCFT, with some UV symmetries decoupling in the IR. We will discuss such exceptional cases in section 4.3. In this case, the affine subspace is one-dimensional and can be represented as
@@ 3.4 Half-index and modular functions
Finally, we consider an ${\cal N}=2$ gauge theory $T[K,\{{\cal O}_{I}\}]$ defined on the Euclidean half-space $\mathbb{C}\times\mathbb{R}_{\leq 0}$. The local operators on the boundary are counted by the half-index, which can be understood as the partition function on the cigar geometry $D^{2}\times S^{1}$, with a choice of boundary condition $\mathbb{B}$ specified on the boundary torus Gadde:2013sca ; Gadde:2013wq ; Yoshida:2014ssa ; Dimofte:2017tpi . Let $(z,\bar{z},t)$ be the local coordinate around a point on the boundary placed at $t=0$. The half index is defined by
@@
where $c\neq 0$ is a complex number. [^11] Under the $D_{c}$ boundary condition, the reference R-charge of the gauge-variant operator $\Phi$ is no longer a free choice in the F-maximization procedure (see footnote 7), since the bulk $U(1)^{r}$ gauge symmetry reduces to the boundary global symmetry $U(1)^{r}_{\partial}$, which are further fully broken for generic $c$. To preserve the $U(1)_{R}$ symmetry, one must choose $R_{*}(\Phi)=0$, which is the choice we adopt (see section 2). As emphasized in Dimofte:2017tpi , this choice should be viewed as part of the definition of the superconformal boundary condition. We denote this boundary condition by $D_{c}$, a deformed Dirichlet boundary condition. The half-index of ${\cal T}[K,\{{\cal O}_{I}\}]$ with the insertion of a supersymmetric Wilson loop $W_{\vec{Q}}$ then reads
@@
with the Pochhammer symbol $(q)_{n}:=\prod_{i=1}^{n}(1-q^{i})$ defined in the usual way. The summation over $m$ labels the boundary monopole operators discussed above. Notice that with our convention (16), the anomaly coefficient of the boundary global symmetry $U(1)^{r}_{\partial}$ (which are fully broken for generic $c$), is identified with the CS level $K$.
@@ 4.1.1 $r=1$
    $K=(2)=C(A_{1})$,  [^14] Taking into account the background CS level 1/2-1/2 in the ${\cal T}_{\Delta}$ theory (2), this theory corresponds to the $U(1)_{3/2}$ theory coupled to a single chiral multiplet of charge +1+1, which we denote simply as $U(1)_{3/2}+\Phi$. This theory is IR dual to $U(1)_{-3/2}+\Phi$ Gang:2018huc . The fact that the Dirichlet half-index coincides with the vacuum character of $M(2,5)$ was first observed in Gang:2023rei .
@@ A.1 General structure of partition functions on $\mathcal{M}_{g,p}$
The effective twisted superpotential $\mathcal{W}(u,v)$ is broken down into two pieces, one originating from the Chern-Simons terms of the theory and one originating from the one-loop contribution of the $\Phi_{i}$ [^20] Importantly this is all done in the $U(1)_{-\frac{1}{2}}$ quantization” which arises due to the parity anomaly that exists in 3D gauge theories as discussed in Closset:2018ghr .: